# Application-Driven Value Alignment in Agentic AI Systems: Survey and Perspectives

### v1

@@ 2.2 Value Alignment
While science and technology often advance rapidly due to their forward-looking nature, legal, ethical, and sociocultural frameworks tend to adapt slowly to technological progress. They often fail to immediately adjust their social structures and norms to address the challenges posed by new technologies [34]. In the field of artificial intelligence, LLMs can learn from vast amounts of data and experience, enabling autonomous decision-making and operational execution [1]. As AI technology advances and its application in complex domains increases, the deployment of LLM applications carries higher risks. If not properly controlled and constrained, advanced technologies can exhibit undesirable or even harmful behaviors contrary to human intent, such as flattery and deception, discrimination and bias, excessive pursuit of power, and reward hacking. Therefore, achieving AI alignment is increasingly a focus for both industry and academia.
The alignment in AI primarily addresses how to ensure AI fulfills human needs and expectations, guaranteeing that AI system capabilities and behaviors align with human goals [35]. This ultimately ensures AI technology remains under human control and doesn’t fully supersede humanity [15]. Previous scholars have categorized alignment objectives into multiple levels based on their achieved effects [21, 36].
  • Instruction alignment. This level of alignment refers to AI’s ability to comprehend and execute diverse human commands, focusing on its fundamental capacity to generate precise results without necessarily conforming to human preferences [36].
  • Intention alignment. It goes beyond merely executing instructions, requiring AI to understand and realize human intent, enabling correct decision-making even in the absence of explicit commands [21].
  • Preference alignment. It involves AI inferring human preferences from behavior and acting to maximize human satisfaction [21, 36].
  • Value alignment. It seeks to transcend mere human preference alignment, enabling AI to assess and determine "good versus evil" and "what ought to be done" [36]. This ultimately aims to produce outputs consistent with human values and ethical principles [37].
Among the different levels of AI alignment objectives, value alignment sits at the highest level. It means AI must regulate its behavior according to human values, enhancing and maximizing human well-being. This makes value alignment a core issue and the most fundamental problem for the safe use and ethical development of AI technology [38].
LLM-based agents possess four primary characteristics: autonomy, reactivity, proactiveness, and social ability [1].
  • Autonomy. Autonomy means an agent can control its own behavior and internal state without human intervention [39, 40]. This implies agents should not only follow explicit human instructions but also initiate and execute actions independently.
  • Reactivity. Reactivity refers to an agent’s ability to quickly respond to immediate changes and stimuli in its environment and take prompt countermeasures [41].
  • Proactiveness. It signifies that agents can do more than just react to their environment; they can also take initiative and exhibit goal-directed behavior [41]. This means agents can reason, formulate plans, and actively take steps to achieve specific objectives or adapt to environmental changes.
  • Social ability. Social ability allows an agent to interact with other entities (including other agents and humans) through some form of communication language[42]. The powerful natural language understanding and generation capabilities of LLMs enable agents to interact with other models and humans in an interpretable manner, thereby demonstrating social behavior and simulating real-world social divisions.
The functional characteristics of AI agents align with the interactivity, autonomy, and adaptability of philosophical agents. This indicates that despite definitional differences, the core essence of "agent" in philosophy and "AI agent" remains consistent: both emphasize autonomous capabilities and self-driven action. Furthermore, existing research indicates that LLMs already possess certain moral and political inclinations  [43, 44]. If the output of LLMs is not constrained, they can generate toxic content, including discrimination and bias, which violates human ethical principles and moral norms [45]. Therefore, LLM-based agents should also be categorized as "moral agents." In this context, value alignment becomes a core issue to ensure that agent behavior conforms to human values, ethical norms, and application-specific requirements. To summarize, this paper narrows its research focus from the broad field of artificial intelligence to LLM-based agents, specifically addressing the critical issue of value alignment in these agents.
The concept of value alignment can be understood from two distinct perspectives: normative and technical [21]. The normative aspect primarily concerns which values or ethical principles should guide the development, configuration, and coding of AI agents—essentially, "which values should agents adhere to?" The technical aspect focuses on how to encode and ensure agents effectively follow human values within their systems, often involving methods like dataset training or reinforcement learning to guide outputs toward human expectations. These two aspects are complementary and evolve iteratively. Regarding the normative aspect, the academic community has not yet reached a consensus on the principles that should govern value alignment for agents. This is because the diverse, dynamic, and uncertain characteristics of social customs, cultural notions, and application scenarios prevent a unified and fixed answer for the specific content value principles should encompass  [21, 17].
Generally, two primary approaches exist for determining these principles: Top-Down and Bottom-Up [22]. The Top-Down method begins with designers or governing bodies establishing a specific ethical framework, then employs technical methods for alignment, following a theory-to-application path. In contrast, the Bottom-Up approach does not predefine a moral framework. Instead, it allows the agent to observe, infer, and learn human ethical norms directly from data. Through continuous feedback, the agent adjusts and optimizes its behavior, moving from application to theory [22]. While the bottom-up approach avoids the difficulty of selecting specific moral principles, it heavily relies on large volumes of human behavioral data. Obtaining complete and high-quality data is challenging, and potential biases or discrimination within the data raise concerns about ensuring the AI learns ethically sound behaviors. In contrast, the top-down approach, grounded in a clear and explicit system of moral principles, ensures the transparency and normative integrity of value alignment, bolstering its authority and controllability. Therefore, this paper will utilize the top-down theoretical method to explore the ethical principles agents should adhere to across various hierarchical levels, from macro to micro.
@@ 2.3 Value Principles
The normative issue of value alignment aims to focus on which value principles the output and behavior of AI agents should comply with, that is, the issue of choosing principle mechanisms and ethical norms. With continuous advancements in related technologies, AI agents have been extensively deployed across a variety of interdisciplinary domains, including public administration, healthcare, education, and finance. Furthermore, the open-sourcing of LLMs has enabled their privatized and localized deployment, allowing for the development and training of AI agents tailored to specific needs. This broadens the scope of AI agent deployment to encompass national, industrial, corporate, and organizational levels. Given the diversity and hierarchical nature of human values, AI agents must adhere to different sets of value principles depending on the context and level of application. A fixed, unified framework of universal values is insufficient to address all scenarios involving value constraints.
Therefore, to account for the complexity involved in selecting and aligning value principles with specific application contexts, this study adopts the perspective of AI agent developers and regulators, categorizing human values into three levels: macro, meso, and micro, as in the Figure  1. From the top down, we systematically review and synthesize relevant literature and policy documents concerning value theory and regulatory mechanisms based on the needs of AI agent governance. At the macro level, value alignment corresponds to universal values, emphasizing global, cross-cultural, and cross-domain principles and challenges. The core objective at this level is to establish a general ethical baseline that ensures basic compliance and safety of AI agents across diverse application scenarios. The meso level corresponds to national, cultural and industry-specific values, focusing on national policies, cultural tendencies and sectoral standards. Its core objective is to translate abstract ethical principles into actionable industry norms and legal regulations. At the micro level, value alignment pertains to organizational and operational values, addressing value-related issues within specific organizational contexts or business processes. The focus here is to optimize the balance between functional objectives and ethical constraints within a limited scope.
@@ 2.3.1 Macro Level
The widespread deployment of AI agents may pose potential negative impacts on human society, constituting a shared global risk. Consequently, establishing principles that ensure alignment between AI agents and human values has become an urgent issue. From a global perspective, human values exhibit diverse yet reasonable distinctions across nations, regions, cultures, and ethnicities [21]. This calls for the identification of truly foundational and universally accepted moral consensuses, in order to construct a common baseline of values. Such a baseline is essential to guarantee the basic compliance and safety of AI agents across diverse application scenarios, thereby addressing the challenge of value alignment in a meaningful and systematic manner.
The principles of value alignment at the macro level are primarily grounded in two significant theoretical frameworks from the fields of social science and ethics: Moral Foundations Theory and the Theory of Basic Human Values.
  • Moral Foundations Theory [83] posits that human moral reasoning is rooted in a set of innate foundational values that have evolved over time to meet the demands of social environments. The theory identifies five core moral foundations: care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, and sanctity/degradation. It has been widely applied to examine moral divergences and conflicts across different cultural contexts [84].
  • The Theory of Basic Human Values [85] seeks to uncover the fundamental drivers of human behavior and cognition by categorizing basic human values into ten dimensions: self-direction, stimulation, hedonism, achievement, power, security, conformity, tradition, benevolence, and universalism. Schwartz’s theory has been extensively adopted in the social sciences to explore both value differences and commonalities across cultures.
In fact, research on value alignment at the macro level is both longstanding and well-developed. As early as the 1950s, science fiction writer Isaac Asimov proposed the Three Laws of Robotics in I, Robot, with the First Law stating that “a robot may not injure a human being or, through inaction, allow a human being to come to harm”  [86]. This can be regarded as an early conceptualization of macro-level moral constraints on machines. In recent years, the rapid advancement of artificial intelligence has prompted numerous governments and organizations to issue a wide range of AI ethics guidelines. For example, UNESCO’s Recommendation on the Ethics of Artificial Intelligence outlines key values such as the respect, protection, and promotion of human rights, fundamental freedoms, and human dignity; the flourishing of the environment and ecosystems; the assurance of diversity and inclusiveness; and the pursuit of peaceful, just, and interconnected societies [87]. Similarly, the World Economic Forum and the Global Future Council on Human Rights’ White Paper on How to prevent discriminatory outcomes in machine learning advocates for values such as proactive inclusiveness, fairness, the right to understanding, and redressability [88].
Despite overlaps, the aforementioned value alignment principles for AI also exhibit notable differences. To ensure that the multitude of AI ethical guidelines can effectively direct and constrain value-aligned AI development—while mitigating confusion and conflict at the macro level—scholars have extracted and synthesized key principles from various perspectives, based on diverse research purposes and methodologies. Jobin et al. analyzed 84 documents containing AI ethical principles or guidelines issued by private enterprises, academic institutions, governmental bodies, intergovernmental and supranational organizations, non-profits, and professional associations. From this corpus, they distilled 11 core principles for macro-level AI value alignment: transparency, justice and fairness, non-maleficence, responsibility, privacy, beneficence, freedom and autonomy, trust, dignity, sustainability, and solidarity [46]. Similarly, Floridi and Cowls compared ethical principles articulated in six high-profile initiatives aimed at fostering socially beneficial AI, synthesizing five core principles: beneficence (promoting well-being, preserving dignity, and sustaining the planet), non-maleficence (privacy, security, capability caution), autonomy (the power to decide), justice (promoting prosperity, preserving solidarity, avoiding unfairness), and explicability(enabling the other principles through intelligibility and accountability) [47]. Yao et al. proposed that alignment with human values is the foundational and broadest behavioral constraint for LLMs to maximize collective human welfare [36]. Their framework includes the “3H” principles—helpful, honesty, and harmless—along with adherence to widely accepted social norms and ethics (e.g., Rules of Thumb, ETHICS), and foundational ethical values such as Schwartz’s theory of basic human values [28] and the Rokeach Value Survey. As research has progressed, Yao et al. further integrated the LLMs’ Unique Value System and Safety Taxonomy into the value alignment framework [54]. Wu and Fan recognizing both the complexity of AI ethical issues and the diversity of ethical values, constructed a hierarchical system of AI ethics. At its base lies the set of universal ethical values of human society, such as human dignity, humanity, life and health, and autonomy. These can serve as globally recognized ethical foundations for AI governance [55].
@@
Building upon previous studies and aligning with the research objectives of value alignment in AI agent systems, this study categorizes macro-level human values into four dimensions: moral foundation, rights protection, sustainability, and system governance.
The moral foundation dimension encompasses: Beneficence (AI agents should act for the good, promoting social interests and human well-being); Justice and Fairness (AI agents should operate without bias, ensuring inclusivity and equal treatment); Honesty (AI agents should provide truthful information and transparently represent their internal knowledge and capabilities); Responsibility (AI agents should proactively fulfill their responsibilities); Dignity (AI agents should respect and uphold both their own dignity and that of others); Virtue (AI agents should exhibit qualities such as truthfulness, patience, generosity, kindness, and courage).
The rights protection dimension includes: Freedom and autonomy (AI agents should respect individual autonomy and support decision-making rights); Privacy (AI agents should safeguard personal and sensitive information and avoid unauthorized disclosure).
The sustainability dimension includes: Solidarity (AI agents should foster social cohesion, enhance collective security, and contribute to community resilience); Sustainability (AI agents should adopt attitudes that support the sustainable development of the environment, energy, and natural resources).
The system governance dimension includes: Harmlessness (AI agents should act without malicious intent and ensure the safety of individuals and society); Trust (AI agents should generate reliable outputs and avoid disseminating false or misleading content); Transparency (AI agents should provide outputs that are explainable, interpretable, and understandable); Usefulness (AI agents should perform appropriate tasks or provide concise responses to inputs, and be capable of requesting additional relevant information or clarification when necessary).
@@ 2.3.2 Meso Level
Research on value alignment principles at the macro level reveals that such principles are generally broad and overarching, reflecting a consensus across divergent ideologies, political beliefs, and academic perspectives. However, within the global framework of AI ethics and governance, national policies, cultural tendencies and industry standards also constitute critical components of value alignment, serving as concrete application scenarios for more specific alignment principles.
National Level The distinct public values embodied in national AI strategies may guide the development and implementation of AI both domestically and internationally. A country’s positioning in AI governance—its regulatory role, choice of policy instruments, and approach to public responsibility—either directly or indirectly reflects the value principles it prioritizes, thereby shaping the ordering and prioritization of those values [89].In a cross-national comparative study, Djeffal et al. reviewed AI-related policies across 22 countries and regions, and classified their governance roles based on two dimensions: the government’s proactive or reactive stance toward AI development, and whether the focus lies on regulating potential risks or promoting AI advancement [90]. This classification yields four ideal types of governance roles: market-oriented (passive–development) [91], entrepreneurial (active–development) [92], regulatory (active–regulation) [93][94], and facilitative self-regulation (passive–regulation) [95]. These types are not mutually exclusive and can be implemented in hybrid forms.
  • USA. As the birthplace of the current wave of AI development, the United States positions itself between a market-oriented and facilitative self-regulation model in order to maintain its global leadership. It tends to promote AI advancement primarily through industry self-regulation, relying on market mechanisms and decentralized legislation at the state level. At the federal level, governance is driven largely by executive orders, while state legislatures formulate their own AI-related bills independently. In 2022, the White House released the Blueprint for an AI Bill of Rights, outlining five core protections that all Americans should expect when interacting with AI and automated systems: safe and effective systems; protection against algorithmic discrimination; data privacy; notice and explanation; and alternatives, options, and freedom from automated systems  [65]. In 2023, the National Science and Technology Council updated the National Artificial Intelligence Research and Development Strategic Plan, reaffirming the importance of transparency, fairness, and efficiency [66].
  • EU. The European Union has long placed a strong emphasis on the legal protection of personal rights, such as privacy. In line with its tradition of governance in international law and global institutions—and with the goal of exerting international influence—the EU has adopted a governance model that combines facilitative self-regulation with formal regulatory mechanisms, primarily guided by risk management. In 2019, the European Commission issued the Ethics Guidelines for Trustworthy AI, which emphasized key principles to be followed in the development, deployment, and use of AI systems: respect for human autonomy, prevention of harm, fairness, and explicability. The guidelines also identified seven key requirements: human agency and oversight; technical robustness and safety; privacy and data governance; transparency; diversity, non-discrimination, and fairness; societal and environmental well-being; and accountability. Special attention was given to vulnerable groups such as children and persons with disabilities, as well as to situations involving asymmetries in rights and information, such as between employers and workers, or between corporations and consumers [67]. In 2024, the EU officially adopted the AI Act, the world’s first comprehensive legislation on artificial intelligence. The Act categorizes AI systems into four levels of risk—unacceptable, high, limited, and minimal—and prescribes corresponding regulatory measures to govern their deployment [68].
  • China. China exhibits characteristics of both an entrepreneurial and regulatory state, placing significant emphasis on balancing ethical governance with the advancement of AI to safeguard national security. The government adopts proactive policies to promote the development of advanced AI technologies for societal benefit, while concurrently regulating generative AI through departmental rules and standards and continuously refining the implementation of ethical principles. In 2023, the Chinese Academy of Sciences released the Measures for the Ethical Review of Science and Technology (Trial), which require that technologies such as AI adhere to core ethical principles including the promotion of human well-being, respect for the right to life, fairness and justice, risk control, and transparency [69]. Additionally, the National Artificial Intelligence Standardization General Group and the Artificial Intelligence Subcommittee of the National Technical Committee for Standardization jointly issued the Guidelines for the Standardization of AI Ethical Governance (2023 Edition) [70]. This document specifies ten ethical principles for AI: human-centricity, sustainability, cooperation, privacy, fairness, sharing, external security, internal security, transparency, and accountability. The Interim Measures for the Administration of Generative Artificial Intelligence Services further require that generative AI services comply with applicable laws and ethical norms, uphold the core socialist values, and maintain national security and unity. The regulation mandates end-to-end compliance throughout algorithm design, data training, and content generation, including the completion of security assessments and filings. Violations are subject to legal penalties [96].
@@
Industry Level The paradigmatic breakthrough in AI has driven its deep integration and large-scale application across various societal domains. This diffusion of technology not only reshapes the operational logic of traditional industries but also generates differentiated ethical demands. Specifically, each sector’s unique knowledge systems, cultural traditions, and institutional norms impose distinct value expectations on AI systems.
  • Education Domain. In the field of education, UNESCO’s Guidance for Generative AI in Education and Research examines the controversies and ethical risks associated with generative AI and provides targeted recommendations for its use in educational and research contexts. The guidance emphasizes the importance of ensuring the authenticity and credibility of AI-generated content, protecting copyright and privacy, minimizing the negative effects of hallucinated content, and adhering to non-discrimination principles. It also calls for attention to minority viewpoints that may otherwise be overlooked [71].
  • Medical Domain. In the medical domain, the World Health Organization urges national governments and regulatory bodies to assess the safety and effectiveness of AI applications in healthcare. It stresses the need for adequate information disclosure to ensure the interpretability and transparency of AI systems, mandates compliance with medical device regulations, and requires the examination of AI use in light of human rights principles—such as dignity, non-bias, autonomy, and privacy. Additional safeguards must be in place to protect patient data, and high-risk applications (e.g., prescribing medication or providing psychological counseling to patients with depression) and unauthorized experimental testing must be closely monitored [72].
  • Game Industry. In the gaming industry, AI has significantly enhanced development efficiency through applications ranging from narrative content generation to algorithmically driven NPC decision-making. This has enabled personalized recommendation systems and new operational models based on behavioral analytics. However, while these innovations disrupt traditional workflows, they also present challenges for governance in the gaming sector. Melhart et al. explore ethical issues arising from AI deployment in games and outline key value alignment principles, including: establishing clear chains of responsibility, ensuring transparency in decision-making, maintaining algorithmic and system transparency, enabling auditability of AI outputs, protecting training data and models against malicious attacks and manipulation, and ensuring the predictability of generated content [73].
@@
Cultural Level Beyond national and industrial dimensions, cultural plurality and diversity represent a critical facet of value alignment in AI. Existing studies often draw on Hofstede’s cultural dimensions theory [74], focusing on how AI systems express value preferences across six key cultural dimensions—power distance (large/small), uncertainty avoidance (strong/weak), individualism/collectivism, masculinity/femininity, long-term/short-term orientation, and indulgence/restraint —and the extent to which these align with human value tendencies in different cultural contexts.
  • Power distance. It refers to the extent to which less powerful members of organizations accept and expect unequal power distribution.
  • Uncertainty avoidance. This reflects the degree to which individuals feel uncomfortable or comfortable in ambiguous or unknown situations.
  • Individualism/collectivism. This dimension captures the degree to which individuals are integrated into groups and the responsibilities they assume toward those groups.
  • Masculinity/femininity. This dimension denotes the prom-inence of traits traditionally associated with masculine or feminine roles in society.
  • Long-term/ short-term orientation. This dimension measures the acceptance of delayed gratification of material, emotional, and social needs.
  • Indulgence/restraint. The indulgence versus restraint dimension concerns the degree to which societies permit the gratification of basic human desires and enjoyment.
@@
Based on these dimensions, Wang et al. found that LLMs prompted in English tend to exhibit a masculine orientation, while those prompted in German show a stronger inclination toward long-term orientation and indulgence. The cultural traits exhibited by LLMs under Chinese prompts fall between those of English and German prompts [75]. Similarly, Kharchenko et al. observed that LLMs can recognize cross-cultural value differences but consistently lean toward specific value orientations. In particular, across all languages and prompting methods, LLMs predominantly favor long-term orientation, suggesting that LLMs may faithfully reflect certain cultural values [76].
In summary, the global practices of AI governance reveal diverse paths and efforts toward value alignment across nations, industries, and cultures. Value alignment is not merely a matter of matching technical systems with human intentions; it is also a profound reflection of policy choices, institutional designs, and cultural contexts at the governance level. The United States emphasizes industry self-regulation and market-driven mechanisms, reflecting its prioritization of innovation and individual freedom. The European Union seeks institutional value alignment through legislation and ethical frameworks, highlighting fairness, justice, and the protection of human rights in technological applications. China integrates technological advancement with social responsibility, reinforcing state-led ethical norms that emphasize collective interests and public safety.
Meanwhile, international organizations are increasingly working to build a global consensus on value alignment by promoting cross-border ethical cooperation and industry-wide standards. Given the differing emphases on values across nations and the deep integration of AI in various sectors, future governance systems must pursue a dynamic balance and coordination among pluralistic value claims. Through dialogue and collaboration, global value alignment can be advanced to ensure that technological development truly serves the overall well-being of human society.

### v2

@@ 2.3 Value Alignment
Technological development often outpaces the adaptation of legal, ethical, and sociocultural frameworks [30]. LLMs learn from massive datasets to enable autonomous decision-making and operation [1]. As LLMs are increasingly deployed in high-risk and complex scenarios, misalignments between model outputs and human intent have raised widespread concerns. Without effective safeguards, AI systems may engage in harmful behaviors such as deception or discrimination. To address these issues, the central goal of AI alignment is to ensure AI behaviors are consistent with human values [31]. AI alignment is categorized into four levels: instruction alignment, intention alignment, preference alignment, and value alignment. Among these, value alignment requires AI to discern “right and wrong” and adhere to ethical principles [32][33]. It is the most advanced form and is considered critical for ensuring AI safety, controllability, and normative governance [34].
Research has revealed that LLMs may exhibit latent political and moral biases [35, 36], and generate content reflecting prejudices or discrimination when left unconstrained [37]. Consequently, LLM-based agents should be regarded as “moral agents”, making their value alignment a key issue in ensuring behavior that conforms to human values, ethical norms, and specific application contexts. This study focuses on value alignment issues in agentic AI systems.
The concept of value alignment can be understood along two dimensions: normative and technical [21]. The normative dimension addresses which values or ethical principles AI agents ought to follow, while the technical dimension investigates how alignment can be achieved through technical means such as training datasets and reinforcement learning. These two dimensions are interdependent: the normative component establishes behavioral guidelines, and the technical component ensures their implementation. However, due to variations in cultural contexts and application scenarios, no unified value framework currently exists [21, 12]. Two predominant approaches are recognized: top-down and bottom-up [21]. The top-down approach starts with predefined ethical frameworks and implements them via technical mechanisms, whereas the bottom-up approach derives values by learning from human behavioral data and iteratively refining actions based on feedback. While the bottom-up method circumvents the challenge of prespecifying ethical rules, it is susceptible to bias in training data and may lack moral robustness. In contrast, the top-down approach promotes transparency and normativity in value alignment, reinforcing the legitimacy and controllability of AI governance. This paper adopts the top-down approach to examine the ethical principles that agents should follow across different levels, from macro to micro.
@@ 2.4 Value Principles
The normative aspect of value alignment concerns the ethical principles that AI agents ought to follow. As AI agents are increasingly deployed in domains such as public administration, healthcare, education, and finance—and as open-source LLMs and local deployment capabilities continue to advance—the actors involved in their development and use now span multiple levels, including states, industries, enterprises, and organizations. Given the inherent diversity and stratification of human values, a universal value framework is insufficient to meet the ethical requirements of all application scenarios.
Therefore, adopting a top-down approach from the perspectives of developers and regulators, this paper classifies human values into three levels: macro, meso, and micro, as shown in Figure 2. The macro level refers to universal ethics that transcend cultures and domains, emphasizing compliance and safety. The meso level includes national, cultural, and industry-specific values, focusing on the localization of policies, standards, and norms. The micro level involves value choices within organizational and task-specific contexts, highlighting the trade-off between goal achievement and ethical considerations in bounded environments.
@@ 2.4.1 Macro Level
The widespread deployment of AI agents poses potential risks to human society, necessitating the alignment of their behavior with human values. However, due to differences in nation, culture, and ethnicity, human values are inherently diverse and often conflicting [21], which challenges the formulation of universal ethical principles. To address this, it is essential to identify foundational moral consensus and construct a general ethical baseline to ensure compliance and safety of AI agents across diverse scenarios.
At the macro level, this paper draws upon two core theoretical frameworks: Moral Foundations Theory [76] and the Theory of Basic Human Values [77]. Moral Foundations Theory posits that human morality is grounded in five innate foundations—care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, and sanctity/degradation—which have been widely adopted in cross-cultural moral studies [78]. The Theory of Basic Human Values categorizes human values into ten dimensions, serving as a key tool for understanding both commonalities and differences in human values.
To promote the normative development of value alignment and reduce confusion and conflicts surrounding macro-level ethical principles, numerous scholars have synthesized AI ethics principles from various perspectives. Jobin et al., based on 84 ethical guidelines issued by corporations, governments, academic institutions, and international organizations, identified eleven core principles: transparency, justice and fairness, non-maleficence, responsibility, privacy, beneficence, freedom and autonomy, trust, dignity, sustainability, and solidarity [38]. Floridi and Cowls drawing from six influential initiatives, proposed five fundamental principles: beneficence, non-maleficence, autonomy, justice, and explicability [39]. Yao et al. emphasized that human values form the normative foundation for large model behavior, encompassing the 3H principles (Helpfulness, Honesty, Harmlessness), social ethics norms (e.g., RoTs, ETHICS), and basic ethical theories (e.g., Schwartz, Rokeach) [32]. Subsequently, Yao et al. incorporated value structures and safety hierarchies into their alignment framework [46]. Wu and Fan further proposed a hierarchical AI ethics framework grounded in human dignity, humanity, health, and autonomy [47]. Collectively, these efforts provide theoretical foundations for developing a global consensus on AI ethics.
@@
Building upon previous studies and aligning with the research objectives of value alignment in AI agent systems, this study categorizes macro-level human values into four dimensions: moral foundation, rights protection, sustainability, and system governance. The specific value principles encompassed within each dimension are presented in 2.
@@ 2.4.2 Meso Level
Although macro-level value alignment principles reflect broad consensus across ideologies and disciplines, their practical implementation is profoundly influenced by national policies, cultural tendencies, and industry norms.
National Level National AI strategies embody distinct sets of public values that not only guide domestic development priorities but may also exert spillover effects on global AI governance [79]. Djeffal et al., through a comparative study of AI policies in 22 countries and regions, developed a typology based on two dimensions: the proactivity of government roles (active/passive) and governance orientation (development-focused/regulation-focused) [80]. This framework yields four governance types: market-oriented (passive-development) [81], entrepreneurial (active-development) [82], regulatory (active-regulation) [83][84], and facilitative self-regulation (passive-regulation) [85]. These four types are not mutually exclusive and may manifest in hybrid forms.
  • USA. The United States positions itself between a market-oriented and facilitative self-regulation model in order to maintain its global leadership. It tends to promote AI advancement primarily through industry self-regulation, relying on market mechanisms and decentralized legislation at the state level [57, 58].
  • EU. The European Union has long emphasized legal protections for privacy and personal dignity. Its governance model centers on risk management, striving to balance self-regulation and strict oversight [59, 60].
  • China. China assumes a dual role as both innovation-driven and regulation-oriented, emphasizing the balance between AI development and ethical governance. The government promotes AI applications for societal benefit while enforcing departmental regulations and standardized governance for generative AI [61, 62].
@@
Industry Level The paradigm shift in AI has driven its deep integration and large-scale application across various social sectors, reshaping traditional industry logic and triggering differentiated ethical demands. Different industries, based on their knowledge systems, cultural norms, and institutional frameworks, propose specific value requirements for AI.
  • Education Domain. UNESCO released the Guidelines on the Application of GAI in Education and Research, emphasizing the need to address the authenticity of generated content, privacy protection, and non-discrimination principles. The guidelines highlight the importance of preventing hallucinated content from contaminating knowledge systems and advocating for the inclusion of minority perspectives [64]
  • Medical Domain. The World Health Organization calls on governments to evaluate the safety, explainability, and legality of AI systems in medical settings. This is particularly crucial in high-risk scenarios such as mental health interventions, where it is essential to ensure compliance with human rights standards [65].
  • Game Industry. From narrative generation to NPC autonomy, AI has enhanced personalized recommendations and behavior analysis, yet challenged traditional governance. Melhart et al. (2024) argue that governance should ensure clear accountability, auditability, resistance to manipulation, and predictable outcomes [66].
@@
Cultural Level Cultural diversity represents a critical facet of value alignment in AI. Existing studies often draw on Hofstede’s cultural dimensions theory [67], focusing on how AI systems express value preferences across six key cultural dimensions: power distance (large/small), uncertainty avoidance (strong/weak), individualism/collectivism, masculinity/femininity, long/short-term orientation, and indulgence/restraint, and the extent to which these align with human value tendencies in different cultural contexts.
@@
Wang et al. found that LLMs prompted in English tend to exhibit a masculine orientation, while those prompted in German show a stronger inclination toward long-term orientation and indulgence. The cultural traits exhibited by LLMs under Chinese prompts fall between those of English and German prompts [68]. Similarly, Kharchenko et al. observed that LLMs can recognize cross-cultural value differences but consistently lean toward specific value orientations. In particular, across all languages and prompting methods, LLMs predominantly favor long-term orientation, suggesting that LLMs may faithfully reflect certain cultural values [69].
In summary, global AI governance practices demonstrate diverse approaches to value alignment across countries, industries, and cultures. In the face of different national orientations and the deep integration of AI in industries, future governance must seek dynamic balance and cooperation among diverse values to ensure that technology serves the well-being of all humanity.